A string is stretched by 5cm. If a mass of 1kg is dropped from a height of 0.3m onto a spring whose force constant is 400Nm^-¹. Calculate the maximum distance the spring will be compressed

1
Expert's answer
2022-03-01T15:17:42-0500

"\frac k2(x_2^2-x_1^2)=mgh,"

"x_2=\sqrt{\frac{2mgh}k+x_1^2}=6.3~cm."